OpenStudy (arrisjay):

Let f(x) = x + 7 and g(x) = x − 4. Find f(x) ⋅ g(x)

OpenStudy (3mar):

the rule is \[(f.g)(x)=f(x).g(x)\] so \[f(X).g(x)=(x+7)(x-4)=x^2-4x+7x-28=x^2+3x-28\]

OpenStudy (arrisjay):

Ahh, thanks!!!

OpenStudy (3mar):

Thanks for the medal!